OpenStudy (anonymous):

5. A plant can either have a smooth seed (p) or a rough seed (q). Rough seeds are recessive and smooth seeds are dominant. In a population of 100 plants, 32 of them have rough seeds. p^2 + 2pq + q^2 = 1 a) What is q^2? (1 point) b) What is q? (1 point) c) What is p? (1 point) d) If there are 19 plants with rough seeds in a population of 100, what is the allele frequency for smooth seeds? (1 point)

OpenStudy (anonymous):

@marcoduuuh omg I haven't seen you in a long time! How's it going! :D I can help you with your question. Do you understand the Hardy-Weinberg equation at all, or are you completely lost?

OpenStudy (anonymous):

Yeah it's been a long time haha! And no, I'm completely lost /:

OpenStudy (anonymous):

Sorry about that, I was dealing with some serious lag! I had to restart the computer. Okay, well, let's get started. Fisrt, you need to find the dominant and recessive alleles. Do you know how to do that?

OpenStudy (anonymous):

I think it's OpenStudy, it's lagging a lot. And yes, the rough seeds are recessive and smooth seeds are dominant.

OpenStudy (anonymous):

Great! Which means that smooth seeds are "p" and rough seeds are "q" The first equation in Hardy-Weinberg is p + q = 1. Since you have 100 plants, solving this will be a piece of cake! :) we could write the expression like this: p + q = 100 We already know q, because those are the recessive plants: p + 32 = 100 If you solve that, you'll find that p is equal to 68. Do you understand how I did that?

OpenStudy (anonymous):

Okay, OS seems to be experiencing lag, so I will just have to move on. Since we figured out p + q = 1 (or, 68 + 32 = 100), we can solve for part a of your question now. By doing p + q = 1 we completed parts b and c already, so I'll let you fill those out. Now we're going to have to use the HW equation itself. \[p^2 + 2pq + q^2 = 1\] Since we already know what the values for p and q are, you could just plug them into the equation and you will get your answer. I would recommend simplifying it into decimal form, though. You can work it out on your calculator. \[.68^2 + 2(.68)(.32) + .32^2 = 1\] Good luck! I hope this helped!

OpenStudy (anonymous):

Wow, thank you so much for all the help and explanation. Sorry I wasn't here, OS started lagging a lot and I couldn't log on to the site. Thank you again for everything, it really helped.

OpenStudy (anonymous):

Oh, you're welcome! Yeah, the lag was terrible. I'm really glad that I was still able to help you, though! :D